# CEASEFIRE: An AI-powered system for combatting illicit firearms trafficking

### v1

@@ Abstract
Modern technologies have led illicit firearms trafficking to partially merge with cybercrime, while simultaneously permitting its off-line aspects to become more sophisticated. Law enforcement officers face difficult challenges that require hi-tech solutions. This article presents a real-world system, powered by advanced Artificial Intelligence, for facilitating them in their everyday work.
1 Introduction
@@
In short, modern technologies have led illicit firearms trafficking to partially merge with advanced cybercrime, while simultaneously permitting its off-line aspects to become more sophisticated as well. Artificial Intelligence (AI) holds the promise of helping LEA practitioners to surpass these challenges, by automating much of the grunt work in which they regularly engage. The adoption of state-of-the-art AI technology and tools by state authorities is very likely to lead to enormous LEA productivity increases, compared to the recent past.
CEASEFIRE (“Advanced versatile artificial intelligence technologies and interconnected cross-sectoral fully-operational national focal points for combating illicit firearms trafficking”) [^1] Project Website: https://ceasefire-project.eu/ tries to achieve exactly that. Funded by the European Commission’s Horizon Europe programme, it is a 3-year innovation project that began in October 2022 and unites the expertise of 7 companies, 5 research organizations, and 9 law enforcement agencies (LEAs) across Europe. The CEASEFIRE project is developing a cutting-edge technical system that integrates innovative AI tools and algorithms to enhance the capabilities of LEAs in combating firearms trafficking and related criminal activities more efficiently. Its objective is to comprehensively and effectively address all of the aforementioned challenges, by leveraging state-of-the-art technology. This article reviews the specific pressing issues that the CEASEFIRE system is designed to face via its use-cases and subsequently provides a broad overview of its technical design and components. These serve as an example of how modern technological advancements can be harnessed to bolster law enforcement efforts and curb the proliferation of illegal firearms.
2 Related Work
@@ 4 The CEASEFIRE System
5 AI for combating firearms trafficking
@@
In general, common technical solutions are utilized for integrating the various system components and providing inter-module communication, such as RESTful Web Services and the respective clients, Apache Kafka for data ingestion, etc. The software assets of the entire system are gathered into the CEASEFIRE Portal Hub, to be easily accessible by the LEA end-users. Its core functionalities revolve around giving access to the CEASEFIRE services and making any produced reports available to the LEA officers. Thus, the Portal Hub offers a comprehensive User Interface where the user, after successfully authenticating, gains an overview of all the CEASEFIRE services that are available to him/her. After selecting the desired service, the user gets redirected to it without the need of repeating the authentication process.
@@
5.1 Cyber-patrolling and cross-border information exchange tools
@@
  • TOR333ttps://www.torproject.org/, a piece of software that facilitates communication with the Dark Web and anonymous navigation in the Surface Web.
@@
5.2 Off-line Trafficking Analysis
Scanned parcel image analysis. This asset is a back-end service that allows automatic detection of firearms components/ammunition in X-ray scans of parcels. It relies on a appropriately pretrained Deep Neural Network (DNN), composed of a Convolutional Neural Network (CNN) backbone and a Vision Transformer head for object detection [7]. The back-end service communicates directly with the respective front-end Web application with a suitable GUI for customs officers. The application pushes to the service requests containing X-ray images and, for each image, receives back the type of object detected and an on-image bounding box (in pixel coordinates). Overall, 10 object classes are supported (e.g., ‘Revolver’, ‘Shotgun’, etc.). After each analysis session, a brief report in the form of a JSON file can be optionally sent to the CEASEFIRE database. This asset powers the CEASEFIRE tool developed for the system’s Use-Case #4.
Firearm type identification. This asset is a back-end service that allows recognition of the model of a firearm in an RGB photograph, taken by a mobile phone at a crime scene. It relies on an appropriately pretrained Vision Transformer DNN for fine-grained visual classification, which incorporates a background subtraction module [8]. The back-end service communicates directly with the respective front-end mobile application that supplies the photographs to be analyzed and receives back the 5 most likely types of weapons per image. After each analysis session, a brief report can be optionally sent to the CEASEFIRE database. This asset powers the CEASEFIRE tools developed for the system’s Use-Case #2.
5.3 On-line Trafficking Analysis
@@
5.4 Information Fusion and cyberphysical traces correlation estimation
@@
6 Supplementary Material
The CEASEFIRE YouTube channel [^8] https://www.youtube.com/@CeasefireProject contains demonstration videos of the project’s front-end applications.
7 Conclusion
@@
8 Acknowledgments
The research leading to these results has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101073876 (Ceasefire).

### v2

@@ Abstract
Modern technologies have enabled illicit firearms trafficking to partially merge with cybercrime, while also allowing its off-line aspects to become increasingly complex. The online trade of firearms, their components, 3D blueprints and illicit substances carried out by criminals on both the surface Web and dark Web is increasingly difficult to address as a consequence of the exponential growth in the amount of information disseminated on the Internet. On the other hand, law enforcement agencies are confronted with significant challenges that require the development of sophisticated technological solutions capable of processing large volumes of data, identifying relevant information in a timely manner and creating networks of connections between potential criminal groups. This article presents a real-world practical system, namely the CEASEFIRE one, powered by advanced artificial intelligence technologies that can assist law enforcement personnel in addressing the above described challenges.
Index Terms:
Combating illicit firearms trafficking, Artificial intelligence (AI), AI integrated system
I Introduction
@@
In short, modern technologies have led illicit firearms trafficking to partially merge with advanced cybercrime, while simultaneously permitting its off-line aspects to become more sophisticated as well. Artificial Intelligence (AI) holds the promise of helping LEA practitioners to surpass these challenges, by automating much of the grunt work in which they regularly engage. The adoption of state-of-the-art AI technology and tools by state authorities is very likely to lead to enormous LEA productivity increases, compared to the recent past. [1]
CEASEFIRE (“Advanced versatile artificial intelligence technologies and interconnected cross-sectoral fully-operational national focal points for combating illicit firearms trafficking”)Project Website: https://ceasefire-project.eu/ tries to achieve exactly that. Funded by the European Commission’s Horizon Europe programme, it is a 3-year innovation project that began in October 2022 and unites the expertise of 7 companies, 5 research organizations, and 9 law enforcement agencies (LEAs) across Europe. The CEASEFIRE project is developing a cutting-edge technical system that integrates innovative AI tools and algorithms to enhance the capabilities of LEAs in combating firearms trafficking and related criminal activities more efficiently. Its objective is to comprehensively and effectively address all of the aforementioned challenges, by leveraging state-of-the-art technology. This article reviews the specific pressing issues that the CEASEFIRE system is designed to face via its use-cases and subsequently provides a broad overview of its technical design and components. These serve as an example of how modern technological advancements can be harnessed to bolster law enforcement efforts and curb the proliferation of illegal firearms.
II Related Work
@@ IV The CEASEFIRE System
V AI for combating firearms trafficking
@@
In general, common technical solutions are utilized for integrating the various system components and providing inter-module communication, such as RESTful Web Services and the respective clients, Apache Kafka for data ingestion, etc. The software assets of the entire system are gathered into the CEASEFIRE Portal Hub, to be easily accessible by the LEA end-users. Its core functionalities revolve around giving access to the CEASEFIRE services and making any produced reports available to the LEA officers. Thus, the Portal Hub offers a comprehensive User Interface where the user, after successfully authenticating, gains an overview of all the CEASEFIRE services that are available to them. After selecting the desired service, the user gets redirected to it without the need of repeating the authentication process.
@@
V-A Cyber-patrolling and cross-border information exchange tools
@@
  • TOR [^2] https://www.torproject.org/, a piece of software that facilitates communication with the Dark Web and anonymous navigation in the Surface Web.
@@
V-B Off-line Trafficking Analysis
Scanned parcel image analysis. This asset is a back-end service that allows automatic detection of firearms components/ammunition in X-ray scans of parcels. It relies on a appropriately pretrained Deep Neural Network (DNN), composed of a Convolutional Neural Network (CNN) backbone and a Vision Transformer head for object detection [8]. The performance of the pipeline is further improved by pretraining the DNN on an unlabeled set of data, using self-supervised training methods [9], particularily useful on low-data settings such as the X-ray illicit firearms domain. The back-end service communicates directly with the respective front-end Web application with a suitable GUI for customs officers. The application pushes to the service requests containing X-ray images and, for each image, receives back the type of object detected and an on-image bounding box (in pixel coordinates). Overall, 10 object classes are supported (e.g., ‘Revolver’, ‘Shotgun’, etc.). After each analysis session, a brief report in the form of a JSON file can be optionally sent to the CEASEFIRE database. This asset powers the CEASEFIRE tool developed for the system’s Use-Case #4.
Firearm type identification. This asset is a back-end service that allows recognition of the model of a firearm in an RGB photograph, taken by a mobile phone at a crime scene. It relies on an appropriately pretrained Vision Transformer DNN for fine-grained visual classification, which incorporates a background subtraction module [10]. The utilisation of diffusion models [11] enables the generation of synthetic images of firearms for use by the DNN in its training phase, thereby enhancing the visual classification rate. The back-end service communicates directly with the respective front-end mobile application that supplies the photographs to be analyzed and receives back the 5 most likely types of weapons per image. After each analysis session, a brief report can be optionally sent to the CEASEFIRE database. This asset powers the CEASEFIRE tools developed for the system’s Use-Case #2.
V-C On-line Trafficking Analysis
@@
V-D Information Fusion and cyberphysical traces correlation estimation
@@
VI Conclusion